Find the exact length of the curve. x = 1 + 12t2, y = 8 + 8Osts 4
Answer #1

y=8+8+ oct 24 x = 1+1 2+² dy Now, 24t, dhe 2 = 244 length of the curre 4 2 dt + dt 0 4 dt Wizada + (29 87? 4 244 vit pa at Le